# MoralCLIP: Contrastive Alignment of Vision-and-Language Representations with Moral Foundations Theory

### v1

@@ Abstract.
Recent advances in vision-language models have enabled rich semantic understanding across modalities. However, these encoding methods lack the ability to interpret or reason about the moral dimensions of content—a crucial aspect of human cognition. In this paper, we address this gap by introducing MoralCLIP, a novel embedding representation method that extends multimodal learning with explicit moral grounding based on Moral Foundations Theory (MFT). Our approach integrates visual and textual moral cues into a unified embedding space, enabling cross-modal moral alignment. MoralCLIP is grounded on the multi-label dataset Social-Moral Image Database (Crone et al., 2018) to identify co-occurring moral foundations in visual content. For MoralCLIP training, we design a moral data augmentation strategy to scale our annotated dataset to 15,000 image-text pairs labeled with MFT-aligned dimensions. Our results demonstrate that explicit moral supervision improves both unimodal and multimodal understanding of moral content, establishing a foundation for morally-aware AI systems capable of recognizing and aligning with human moral values. [^1] All implementations and datasets will be made available after publication.
MoralCLIP, CLIP, Moral, AI, Moral foundations, MFT.
@@ 1. Introduction
Images are among the most powerful stimuli humans encounter, often surpassing text in their ability to instantly convey meaning and evoke emotions (Bavel and Robertson, 2024; Whitehouse et al., 2006). Unlike text, which generally requires greater cognitive effort to interpret, images can evoke intuitive moral responses almost instantly (Harenski et al., 2010; Crone et al., 2018; Keib et al., 2018), with a single image being capable of sparking social movements, changing public opinion, and influencing moral perceptions (Bavel and Robertson, 2024; Keib et al., 2018).
Human communication naturally integrates these modalities to construct meaning (Wong, 2019; Kress, 2010), easily intertwining visual and language modalities. We process the world through this multimodal lens, where text and images interact to create richer, more nuanced understandings (Cheema et al., 2023; Li et al., 2020), particularly in contexts where visual cues might reinforce, contradict, or complicate textual narratives (Keib et al., 2018; Crone et al., 2018; Tamborini et al., 2017). This integration of modalities is especially significant in how people engage with moral content, where the combination of text and images can create more powerful ethical impressions than either medium alone (Cheema et al., 2023; Crone et al., 2018; McCurrie et al., 2018). Recent advances in artificial intelligence have begun to reflect this multimodal integration, with vision-language models like CLIP (Radford et al., 2021), SigLIP (Zhai et al., 2023; Tschannen et al., 2025), and multimodal foundation models such as LLaVA (Liu et al., 2023) and BLIP (Li et al., 2022) bridging the gap between visual and textual information. Although these models have consistently showcased remarkable capabilities in semantic understanding across text and images—enabling applications ranging from visual question answering to cross-modal retrieval and generation (Liu et al., 2023; Li et al., 2022; Dai et al., 2023; Radford et al., 2021; Tschannen et al., 2025; Zhai et al., 2023)—they have not been designed to model or interpret the moral dimensions of content. This represents a significant limitation, as moral evaluations constitute a fundamental aspect of how humans process multimodal information (Harenski et al., 2010; Crone et al., 2018; Keib et al., 2018).
@@
To analyze and understand different expressions of morality, the Moral Foundations Theory (MFT) (Graham et al., 2013) provides a widely adopted framework for evaluating moral judgments across cultures. MFT postulates that moral reasoning is shaped by five innate moral foundations, Table 1Care, Fairness, Authority, In-group, and Purity—which are thought to be universal across human societies. While this theory has been extensively applied to text analysis through both lexicon-based approaches (Hopp et al., 2021; Graham et al., 2009; Matsuo et al., 2019; Zúquete et al., 2023) and language models (Guo et al., 2023; Preniqi et al., 2024; Zangari et al., 2025; Nguyen et al., 2024; Park et al., 2024), recent attempts to extend it to visual analysis have been limited in scope and depth. Thus, a truly integrated multimodal approach to moral foundation analysis remains underdeveloped.
@@
Current methods for moral analysis operate in isolated modalities, with text-only models analyzing written content and image-based approaches relying primarily on textual supervision (Jeong et al., 2022). This uni-modal framing is reflected not only in model architectures but also at the data-level, given most datasets constructed for moral analysis—such as Moral Foundations Twitter Corpus (MFTC) (Hoover et al., 2020), Moral Foundations Reddit Corpus (MFRC) (Trager et al., 2022), MOKA (Zhang et al., 2024), and E2MoCase (Greco et al., 2024)—are exclusively textual, limiting the development of models capable of integrating multimodal moral content. While a few datasets have begun to explore morality in visual content, such as the Social-Moral Image Database (Crone et al., 2018) and the Moral Affective Film Set (McCurrie et al., 2018), these again represent single-modality efforts, offering no way to study how text and images jointly influence moral interpretation. Simultaneously, most vision-based datasets related to values or judgment concentrate on general notions of safety, appropriateness, and societal biases (Hu et al., 2024; Janghorbani and De Melo, 2023; Liu et al., 2025; Seth et al., 2023), rather than pluralistic morality grounded in frameworks like MFT. This is yet another factor that constrains the development of models capable of analyzing or aligning moral information across modalities.
These limitations in model design and dataset construction result in approaches that often fail to capture how moral judgments arise from the interaction between visual and textual elements in real-world scenarios. Furthermore, existing moral analysis methods often reduce morality to a binary concept, without considering the interplay between the different moral foundations (Jeong et al., 2022; Preniqi et al., 2024; Nguyen et al., 2024), effectively overlooking the pluralistic nature of moral reasoning.
To address these challenges, we propose MoralCLIP, a framework that extends vision and language models to capture moral dimensions across both visual and textual modalities. To the best of our knowledge, this is the first attempt to leverage the MFT to create a morally-grounded multimodal embedding space capable of identifying moral content within each modality. Specifically, our contributions are:
@@
  • Morally-Grounded Multimodal Data Augmentation: Leveraged by the expert annotated dataset SMID (Crone et al., 2018), we construct a dataset of 15,000 image-text pairs annotated with MFT-aligned moral labels. The data augmentation process is achieved with a high-precision moral image classifier, Visual Moral Compass, and the generation of short, descriptive captions. We apply this process to the ImageNet (Russakovsky et al., 2015) and LAION-400M (Schuhmann et al., 2021) datasets, while retaining SMID’s-(Crone et al., 2018) expert labels, resulting in the first dataset that explicitly connects moral foundations across both modalities.
While current V&L models, e.g. CLIP (Radford et al., 2021) and SigLIP (Dosovitskiy et al., 2021), excel at semantic understanding, experimental results demonstrate that MoralCLIP successfully captures moral dimensions across modalities, representing a significant first step towards responsible, ethically-aligned multimodal AI that understands not just what we communicate, but the values behind it.
@@ 2. Related Work
Moral Foundations Theory.    Moral Foundations Theory (MFT) is a moral psychology framework for understanding how universal moral foundations shape moral reasoning across diverse cultures. While these foundations are broadly shared across human societies, MFT emphasizes that their specific expressions are deeply influenced by socio-cultural contexts and individual experiences (Graham et al., 2013; Van Den Berg and Corrias, 2023). At its core, MFT posits that moral judgments and decisions arise from emotional, innate evaluations known as moral intuitions (Graham et al., 2013; Tamborini et al., 2017). These intuitions allow individuals to make rapid, unconscious, moral assessments—such as approving or disapproving of an action—based on their moral values.
Specifically, MFT identifies five distinct, yet interconnected moral foundations: Care, Fairness, In-group (or Loyalty), Authority (or Respect), and Purity (or Sanctity). Each of these foundations is structured as a duality, encompassing both a virtue, representing morally positive behavior, and a vice, representing morally reprehensible actions (Graham et al., 2013) (Table 1). In this study, we leverage these five moral foundations as the theoretical backbone of our morally aligned embedding space.
@@
In response to these challenges, recent approaches employed embedding-based methods to model moral similarity, capturing the pluralistic nature of morality (Park et al., 2024), while others have framed moral inference explicitly as a classification task (Preniqi et al., 2024; Guo et al., 2023; Nguyen et al., 2024). For domain invariance, Guo et al. (2023) and Preniqi et al. (2024) employ adversarial training with BERT-based models (Devlin et al., 2019) for binary classification, while  Nguyen et al. (2024) fine-tunes RoBERTa models (Liu et al., 2019) for classification at the foundation level. However, these approaches rely on binary classification, which fundamentally contradicts one of the core principles of the MFT: multiple foundations often co-occur and influence one another. Capturing this complexity requires multi-label models that go beyond isolated, foundation-by-foundation inference.
@@
Morality Encoded in Images.    Few efforts have extended moral analysis to visual content, and even fewer in a truly multimodal fashion (Jeong et al., 2022; Yan et al., 2024). Unlike language, where moral cues are often explicit, visual moral analysis is inherently more complex as images convey morality through visual elements such as expressions, actions, and spatial relationship (Keib et al., 2018; Yan et al., 2024)Jeong et al. (2022) perform zero-shot binary moral classification using CLIP embeddings trained on text-only data (Hendrycks et al., 2021) and apply the resulting classifier to image embeddings. However, this method cannot capture MFT’s dimensional complexity or implicit visual cues. This highlights the need for approaches that can capture MFT’s dimensional complexity through direct visual supervision and true multimodal reasoning.
@@
Vision and Language Pretrained Models.    Large-scale pretraining of vision and language modalities has resulted in significant advancement in cross-modal representation learning. Models such as CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021), SigLIP (Zhai et al., 2023; Tschannen et al., 2025) use dual-encoder architectures to learn aligned embeddings of image-text pairs, enabling strong performance in tasks like zero-shot image classification and image-text retrieval. These models process visual and textual inputs independently, projecting them into a shared embedding space, and have been fine-tuned for various downstream tasks requiring cross-modal understanding (Tschannen et al., 2025; Zhou et al., 2024; Eslami et al., 2023). Recent work has explored modifying these embedding spaces for safety purposes. Safe-CLIP (Poppi et al., 2024) fine-tunes CLIP to reduce sensitivity to NSFW content, redirecting inappropriate inputs to safer embedding regions while preserving CLIP’s original embedding space structure. However, this approach focuses primarily on safety filtering, making it counterproductive for moral classification. In contrast, we leverage the multimodal capabilities of vision and language models to design a morality-aware embedding space that captures the moral dimensions embedded in visual and textual content. Such a space would enable moral analysis grounded in both textual and visual modalities.
@@ 3. A Multimodal Moral Embedding Space
In this section, we propose MoralCLIP, a model that extends the standard CLIP architecture to jointly align image and text representations while incorporating moral information. We considered a morally annotated dataset $\mathcal{D}=\{(v_{1},t_{1},m_{1}),\ldots,(v_{i},t_{i},m_{i}),\ldots\}$, where each tuple is composed of an image $v_{i}$, the corresponding text caption $t_{i}$ and a set of moral labels $m_{i}$ indicating virtue, neutral or vice for each one of the 5 moral foundations, Table 1.
@@ 3.1. CLIP with Implicit Moral
Building upon a pretrained CLIP model, our approach integrates MFT labels into the learning objective, encouraging the model to learn morally grounded embeddings across modalities. As a first approach, we trained CLIP in an implicit setting, using its original contrastive loss objective, with the augmented moral dataset (Section 3.3). We introduce moral information through morally charged images and their corresponding captions, i.e. image-text pairs from dataset $\mathcal{D}$, excluding the moral annotations completely. No architectural or loss modifications are made. This allows us to assess whether CLIP can passively acquire moral information without being guided by the labels.
@@ 3.2. MoralCLIP
(1) $\mathcal{L}_{\text{Total}}=(1-\lambda)\cdot\mathcal{L}_{\text{CLIP}}+\lambda\cdot\mathcal{L}_{\text{Moral}}$
where $\lambda$ controls the influence of the moral alignment component. The moral term $\mathcal{L}_{\text{Moral}}$ penalizes discrepancies between the moral similarity of sample pairs and their semantic similarity in the joint embedding space, computed as the mean squared error:
(2) $\mathcal{L}_{\text{Moral}}=\frac{1}{N}\sum_{i}^{N}\left(\text{sim}(v_{i}^{e},t_{i}^{e})-\text{sim}_{\text{Moral}}(M_{v_{i}},M_{t_{i}})\right)^{2}$
where $\text{sim}(v_{i}^{e},t_{i}^{e})=\langle\hat{v}_{i}^{e},\hat{t}_{i}^{e}\rangle/\tau$ is the scaled cosine similarity between normalized image ($\hat{v}_{i}^{e}$) and text ($\hat{t}_{i}^{e}$) embeddings. Following CLIP’s approach (Radford et al., 2021), we apply temperature scaling $\tau$ to embedding similarity, encouraging more discriminative representations. The moral similarity $\text{sim}_{\text{Moral}}(M_{v_{i}},M_{t_{i}})$ is computed as the scaled Jaccard Index between MFT labels of the image ($M_{v_{i}}$) and text ($M_{t_{i}}$) embeddings:
(3) $\text{sim}_{\text{Moral}}(M_{v_{i}},M_{t_{i}})=2\frac{|M_{v_{i}}\cap M_{t_{i}}|}{|M_{v_{i}}\cup M_{t_{i}}|}-1$
This formulation preserves CLIP’s semantic similarity loss while adding the constraint that semantic similarity should align with moral overlap between samples. This is particularly suitable for our multi-label setting, where samples can be associated with multiple moral foundations simultaneously. To avoid trivial self-similarity effects, we exclude diagonal terms from the loss computation.

### v2

@@ Abstract.
Recent advances in vision-language models have enabled rich semantic understanding across modalities. However, these encoding methods lack the ability to interpret or reason about the moral dimensions of content—a crucial aspect of human cognition. In this paper, we address this gap by introducing MoralCLIP, a novel embedding representation method that extends multimodal learning with explicit moral grounding based on Moral Foundations Theory (MFT). Our approach integrates visual and textual moral cues into a unified embedding space, enabling cross-modal moral alignment. MoralCLIP is grounded on the multi-label dataset Social-Moral Image Database to identify co-occurring moral foundations in visual content. For MoralCLIP training, we design a moral data augmentation strategy to scale our annotated dataset to 15,000 image-text pairs labeled with MFT-aligned dimensions. Our results demonstrate that explicit moral supervision improves both unimodal and multimodal understanding of moral content, establishing a foundation for morally-aware AI systems capable of recognizing and aligning with human moral values. [^1] This paper was published in the Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), Dublin, Ireland. Association for Computing Machinery, New York, NY, USA, 12399–12408. https://doi.org/10.1145/3746027.3758166.
All supplementary material and resources available at https://anaacondez.github.io/moralclip/.
MoralCLIP, CLIP, Moral, Ethics, AI, Moral foundations, MFT.
$c$s: Computing methodologies Computer visionccs: Computing methodologies Natural language processing
@@ 1. Introduction
Images are among the most powerful stimuli humans encounter, often surpassing text in their ability to instantly convey meaning and evoke emotions (Bavel and Robertson, 2024; Whitehouse et al., 2006). Unlike text, which generally requires greater cognitive effort to interpret, images can evoke intuitive moral responses almost instantly (Harenski et al., 2010; Crone et al., 2018; Keib et al., 2018), with a single image being capable of sparking social movements, changing public opinion, and influencing moral perceptions (Bavel and Robertson, 2024; Keib et al., 2018). Human communication naturally integrates these modalities to construct meaning (Wong, 2019; Kress, 2010), easily intertwining visual and language modalities. We process the world through this multimodal lens, where text and images interact to create richer, more nuanced understandings (Cheema et al., 2023; Li et al., 2020). This is particularly evident in moral contexts where visual cues might reinforce, contradict, or complicate textual narratives to create more powerful ethical impressions than either medium alone (Cheema et al., 2023; Crone et al., 2018; McCurrie et al., 2018; Keib et al., 2018; Tamborini et al., 2017). Recent advances in artificial intelligence have begun to reflect this multimodal integration, with vision-language models like CLIP (Radford et al., 2021) and SigLIP (Zhai et al., 2023; Tschannen et al., 2025) bridging the gap between visual and textual information. Although these models excel at semantic understanding across modalities, they have not been designed to model or interpret the moral dimensions of content.
@@
To analyze and understand different expressions of morality, Moral Foundations Theory (MFT) (Graham et al., 2013) provides a widely adopted framework for evaluating moral judgments across cultures. MFT postulates that moral reasoning is shaped by five innate moral foundations, which are thought to be universal across human societies. While this theory has been extensively applied to text analysis through both lexicon-based approaches (Hopp et al., 2021; Graham et al., 2009; Matsuo et al., 2019; Zúquete et al., 2023) and language models (Guo et al., 2023; Preniqi et al., 2024; Zangari et al., 2025; Nguyen et al., 2024; Park et al., 2024), recent attempts to extend it to visual analysis have been limited in scope and depth. Thus, a truly integrated multimodal approach to moral foundation analysis remains underexplored.
@@
Current methods for moral analysis operate in isolated modalities, with text-only models analyzing written content and image-based approaches relying primarily on textual supervision (Jeong et al., 2022). This unimodal framing is reflected not only in model architectures but also at the data level, given most datasets constructed for moral analysis—such as Moral Foundations Twitter Corpus (MFTC) (Hoover et al., 2020), Moral Foundations Reddit Corpus (MFRC) (Trager et al., 2022), Moral Events (Zhang et al., 2024), and E2MoCase (Greco et al., 2024)—are exclusively textual, limiting the development of models capable of integrating multimodal moral content. While a few datasets have begun to explore morality in visual content, such as the Social-Moral Image Database (Crone et al., 2018) and the Moral Affective Film Set (McCurrie et al., 2018), these again represent single-modality efforts, offering no way to study how text and images jointly influence moral interpretation. Simultaneously, most vision-based datasets related to values or judgment concentrate on general notions of safety, appropriateness, and societal biases (Hu et al., 2024; Janghorbani and De Melo, 2023; Liu et al., 2025; Seth et al., 2023), rather than pluralistic morality grounded in frameworks like MFT. Furthermore, existing moral analysis methods often reduce morality to a binary concept, without considering the interplay between the different moral foundations (Jeong et al., 2022; Preniqi et al., 2024; Nguyen et al., 2024), effectively overlooking the pluralistic nature of moral reasoning. To address these challenges, we propose MoralCLIP, a framework that extends vision-language models to capture moral dimensions across both visual and textual modalities. To the best of our knowledge, this is the first attempt to leverage MFT to create a morally-grounded multimodal embedding space capable of identifying moral content within each modality. Specifically, our contributions are:
@@
  • Morally-Grounded Multimodal Data Augmentation: Leveraging the expert annotated dataset SMID (Crone et al., 2018), we construct a dataset of 15,000 image-text pairs annotated with MFT-aligned moral labels. The data augmentation process is achieved with a high-precision moral image classifier, Visual Moral Compass, and the generation of short, descriptive captions. We apply this process to the ImageNet (Russakovsky et al., 2015) and LAION-400M (Schuhmann et al., 2021) datasets, while retaining SMID’s (Crone et al., 2018) expert labels, resulting in the first dataset that explicitly connects moral foundations across both modalities.
While current V&L models, e.g. CLIP (Radford et al., 2021) and SigLIP (Zhai et al., 2023), excel at semantic understanding, experimental results demonstrate that MoralCLIP successfully captures moral dimensions across modalities, representing a significant first step towards responsible, ethically-aligned multimodal AI that understands not just what we communicate, but the values behind it.
@@ 2. Related Work
Moral Foundations Theory.   Moral Foundations Theory (MFT) is a moral psychology framework for understanding how moral foundations shape moral reasoning across diverse cultures. While universal, MFT emphasizes that their specific expressions are deeply influenced by socio-cultural contexts and individual experiences (Graham et al., 2013; Van Den Berg and Corrias, 2023). At its core, MFT posits that moral judgments and decisions arise from emotional, innate evaluations known as moral intuitions (Graham et al., 2013; Tamborini et al., 2017). These intuitions allow individuals to make rapid, unconscious moral assessments—such as approving or disapproving of an action—based on their moral values. Specifically, MFT identifies five distinct, yet interconnected moral foundations: Care, Fairness, In-group (or Loyalty), Authority (or Respect), and Purity (or Sanctity). Each of these foundations is structured as a duality, encompassing both a virtue, representing morally positive behavior, and a vice, representing morally reprehensible actions (Graham et al., 2013) (Table 1). In this study, we leverage these five moral foundations as the theoretical backbone of our morally aligned embedding space.
@@
In response to these challenges, recent approaches employed embedding-based methods to model moral similarity, capturing the pluralistic nature of morality (Park et al., 2024), while others have framed moral inference explicitly as a classification task (Preniqi et al., 2024; Guo et al., 2023; Nguyen et al., 2024). For domain invariance, Guo et al. (2023) and Preniqi et al. (2024) use adversarially trained BERT (Devlin et al., 2019) models, while Nguyen et al. (2024) fine-tunes RoBERTa (Liu et al., 2019) for foundation-level classification. However, these approaches rely on binary classification, which fundamentally contradicts one of the core principles of MFT: multiple foundations often co-occur and influence one another. Capturing this complexity requires multi-label models that go beyond isolated, foundation-by-foundation inference.
@@
Morality Encoded in Images.   Few efforts have extended moral analysis to visual content (Jeong et al., 2022; Yan et al., 2024), where moral cues are conveyed through expressions, actions, and spatial relationships rather than explicit language (Keib et al., 2018; Yan et al., 2024). Jeong et al. (2022) perform zero-shot binary moral classification using CLIP embeddings trained on text-only data (Hendrycks et al., 2021) and apply the resulting classifier to image embeddings. However, this method cannot capture MFT’s dimensional complexity or implicit visual cues. This highlights the need for approaches that can capture MFT’s dimensional complexity through direct visual supervision and true multimodal reasoning.
@@
Vision-Language Pretrained Models.   Large-scale pretraining on vision–language data has driven progress in cross-modal representation learning. Models such as CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021), and SigLIP (Zhai et al., 2023; Tschannen et al., 2025) learn aligned image-text embeddings via dual-encoder architectures, enabling strong performance in zero-shot image classification and image-text retrieval. These models process visual and textual inputs independently, projecting them into a shared embedding space, and can be fine-tuned for various downstream tasks requiring cross-modal understanding (Tschannen et al., 2025; Zhou et al., 2024; Eslami et al., 2023). Recent work has explored modifying these embedding spaces for safety purposes. Safe-CLIP (Poppi et al., 2024) fine-tunes CLIP to reduce sensitivity to NSFW content, redirecting inappropriate inputs to safer embedding regions while preserving CLIP’s embedding space structure. However, this approach targets safety filtering rather than moral understanding. In contrast, we leverage the multimodal capabilities of vision-language models to design a morality-aware embedding space that captures the moral dimensions embedded in visual and textual content. Such a space would enable moral analysis grounded in both textual and visual modalities.
@@ 3. A Multimodal Moral Embedding Space
In this section, we propose MoralCLIP, a model that extends the standard CLIP architecture to jointly align image and text representations while incorporating moral information. We consider a morally annotated dataset $\mathcal{D}=\{(v_{1},t_{1},m_{1}),\ldots,(v_{i},t_{i},m_{i}),\ldots\}$, where each tuple is composed of an image $v_{i}$, the corresponding text caption $t_{i}$, and a set of moral labels $m_{i}$ indicating virtue, neutral, or vice for each of the five moral foundations, Table 1.
@@ 3.1. CLIP with Implicit Moral
Building upon a pretrained CLIP model, our approach integrates MFT labels into the learning objective, encouraging the model to learn morally grounded embeddings across modalities. Initially, we trained CLIP in an implicit setting, using its original contrastive loss objective, with the augmented moral dataset (Section 3.3). We introduce moral information through morally charged images and their corresponding captions, i.e. image-text pairs from dataset $\mathcal{D}$, excluding the moral annotations completely. No architectural or loss modifications are made. This allows us to assess whether CLIP can passively acquire moral information without being guided by the labels.
@@ 3.2. MoralCLIP
(1) $\mathcal{L}_{\text{Total}}=\mathcal{L}_{\text{CLIP}}+\lambda\cdot\mathcal{L}_{\text{Moral}},$
where $\lambda$ controls the influence of the moral alignment component. The moral term $\mathcal{L}_{\text{Moral}}$ penalizes discrepancies between the moral similarity of all sample pairs and their semantic similarity in the joint embedding space, computed as the mean squared error:
(2) $\mathcal{L}_{\text{Moral}}=\frac{1}{N}\sum_{i,j\in B,i\neq j}\left(\text{sim}(v_{i}^{e},t_{j}^{e})-\text{sim}_{\text{Moral}}(M_{v_{i}},M_{t_{j}})\right)^{2}$
where $N$ represents the total number of non-diagonal pairs in the batch, $\text{sim}(v_{i}^{e},t_{j}^{e})=\langle\hat{v}_{i}^{e},\hat{t}_{j}^{e}\rangle/\tau$ is the scaled cosine similarity between normalized image ($\hat{v}_{i}^{e}$) and text ($\hat{t}_{j}^{e}$) embeddings. Following CLIP’s approach (Radford et al., 2021), we apply temperature scaling $\tau$ to embedding similarity, encouraging more discriminative representations. Note that our moral loss does not optimize semantic similarity independently. Rather, it constrains the embedding space so that semantic relationships align with the moral similarity patterns defined by the moral labels. The moral similarity $\text{sim}_{\text{Moral}}(M_{v_{i}},M_{t_{j}})$ is computed as the scaled Jaccard Index between MFT labels of the image ($M_{v_{i}}$) and text ($M_{t_{j}}$) embeddings:
(3) $\text{sim}_{\text{Moral}}(M_{v_{i}},M_{t_{j}})=2\frac{|M_{v_{i}}\cap M_{t_{j}}|}{|M_{v_{i}}\cup M_{t_{j}}|}-1$
This formulation preserves CLIP’s semantic similarity loss while adding the constraint that semantic similarity should align with moral overlap between samples. This is particularly suitable for our multi-label setting, where samples can be associated with multiple moral foundations simultaneously. To avoid trivial self-similarity effects, we exclude diagonal terms from the loss computation. Full training configuration details are provided in Appendix D.